OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, -1) and (-2, -5)? A. 4/5 B. -5/4 C. -4/5 D. 5/4

OpenStudy (smilie2015):

use the equation m = slope =y2-y1/x2-x1

OpenStudy (tanner23456):

slope = (y2-y1)/(x2-x1) (-5--1)/((-2-3) -4/-5 = 4/5 A

OpenStudy (smilie2015):

HINT: (3,-1) (x1,y1)

OpenStudy (smilie2015):

@tanner23456 it is not allowed to give direct answers :)